OpenStudy (anonymous):

A hairdresser is deciding where to open her own studio. if the hairdresser chooses location A, she will pay $1200 per month in rent and will charge $45 per haircut. if she chooses location B, she will pay $1800 per month in rent and will charge $6o per haircut. how many haircuts would she have to give in one month to make the same profit at either location?

OpenStudy (texaschic101):

1200 + 45h = 1800 + 60h @ganeshie8 .....am I setting this up wrong....because this is gonna come out a negative number

OpenStudy (texaschic101):

@jdoe0001 ...can you help

OpenStudy (texaschic101):

I know you got to set them up equal to each other then solve for h, but unless this question is written wrong....I can't get a positive number

OpenStudy (texaschic101):

@e.mccormick ...can you look at this please :)

OpenStudy (jdoe0001):

hmmm

OpenStudy (texaschic101):

am I doing something wrong ?

OpenStudy (jdoe0001):

well... from A = 45x - 1200 B = 60x - 1800 break-even point will be at revenue and expenses cancel each other, so 0 0 = 45x - 1200 0 = 60x - 1800 .. I also get a negative number :|

OpenStudy (texaschic101):

I am beginning to think the problem is written wrong

OpenStudy (jdoe0001):

often times is usually the semantics in the wording

OpenStudy (texaschic101):

I guess your right....well, I tried

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

man I'm a dork... .dunno why I didn't notice.... anyhow the setup is correct

OpenStudy (jdoe0001):

@texaschic101 profit of "y" for each will be y = 45x - 1200 y = 60x - 1800 same profit for both will be 45x-1200 = 60x - 1800 gives a positive value

OpenStudy (texaschic101):

oh...I wasn't subtracting the rent, I was adding it I feel stupid now....thanks for clearing that up